It's me again, I don't know how to solve the system of equations by substitution. {5x-2y= -1
{x+4y= 35
last question then I will leave you alone until next term in school.
Solve the inequality. Give the solution in interval form.

3(5x-1)-2x>13=7
Thank-you so much for your time and patience.

Answer by rapaljer(4551)

To solve by substitution method, you need to solve for one variable in terms of the other in one of the equations. In your problem, the easiest one to solve for is to solve for x in the second equation:

x=-4y +35

Now, substitute this into the first equation:
5x-2y=-1
5(______)-2y=-1
5(-4y+35)-2y=-1
-20y + 175-2y=-1
-22y+175=-1
-22y=-176
y=8

Now, substitute this value y=8 back into they x= equation:
x=-4y+35
x=-4(8)+35
x=-32+35
x=3

The solution is (3,8).

Check:
Substitute both values of x and y into the OTHER equation:
5x-2y=-1
5(3)-2(8)=-1
15-16=-1 , IT CHECKS!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus